# AuToMATo: An Out-Of-The-Box Persistence-Based Clustering Algorithm

### v2

@@ 2.3 The Bottleneck Bootstrap
$\widehat{q}_{\alpha}:=\inf\{z\mid\widetilde{F}(z)\geq 1-\alpha\}$
[…]
2.4 Clustering
Clustering is a fundamental technique in unsupervised learning and thus there is a vast body of literature both on the theoretical foundation as well as on various algorithms and applications of clustering [And73, DHS00]. Very roughly, most clustering algorithms fall into one of four categories: centroid-based algorithms, model-based algorithms, hierarchical clustering algorithms, and density-based algorithms.
The most famous algorithm in the first category is $k$-means, see [Boc07] for a historic overview. The $k$-means algorithm and many other centroid-based algorithms take the number of clusters as an input, rendering them irrelevant for many applications such as Mapper. We thus do not compare our new algorithm against $k$-means and similar algorithms. Similarly, we omit comparisons with model-based algorithms since they assume that the data follows some distribution, which is an assumption we do not make.
Hierarchical clustering algorithms start by considering each data point as its own cluster and iteratively join the two currently closest clusters into a single cluster, stopping either when a certain number of clusters is reached or when any two clusters are far enough apart. As such, they use one of the following two as a hyperparameter: the number of clusters or a distance threshold. Depending on how the distance between clusters is defined, one obtains different hierarchical clustering algorithms. We compare our algorithms against several such algorithms. In single linkage clustering the distance between two clusters $A$ and $B$ is defined as $\min_{a\in A,b\in B}d(a,b)$. Replacing the $\min$ by $\max$, we get the distance function for complete linkage clustering. For average linkage clustering, the distance between two clusters $A$ and $B$ is the average distance between the points, that is, $\frac{1}{|A|\cdot|B|}\sum_{a\in A}\sum_{b\in B}d(a,b)$. Finally, in Ward linkage clustering, the distance between two clusters $A$ and $B$ is defined in terms of their centroids $\mu_{A}$ and $\mu_{B}$ by $\frac{|A|\cdot|B|}{|A\cup B|}\lVert\mu_{a}-\mu_{b}\rVert^{2}$. For surveys on different hierarchical clustering algorithms and a comparison between them we refer to [MT13, SB+19].
Among the most widely used density-based clustering algorithms are DBSCAN [EKSX96] and its adaptation HDBSCAN [CMS13]. DBSCAN has two hyperparameters, a natural number $\mathrm{minPts}$ and a distance threshold $\varepsilon$. Any data point that contains at least $\mathrm{minPts}$ in its ball of radius $\varepsilon$ is called a core point. DBSCAN then assigns two core points to the same cluster whenever the distance between them is at most $\varepsilon$ and also adds all points in their $\varepsilon$-balls to this cluster. Any (non-core points) that are not assigned to any cluster are labelled as noise. HDBSCAN is an extension of DBSCAN that, roughly speaking, performs the same procedure for a range of values for $\varepsilon$ and computes for each potential cluster a stability value, selecting as final clusters those whose stability is greater than the sum of stability values of its subclusters. We include both DBSCAN and HDBSCAN in our comparison.
A clustering method is parameter-free if it does not depend on a good choice of hyperparameters. Several such methods have been introduced in the literature, often for specialized tasks. There is also some recent work on parameter-free clustering methods that are designed for general settings. Among those we mention some for which there are publicly available implementations, which allows us to compare our new algorithm against them. These are FINCH [SSS19] and the persistence-based clustering algorithm [CSTL21] implemented in the Topology ToolKit [CSTL21]. In a nutshell, FINCH proceeds as follows. The algorithm defines a graph on the data points, connecting two points $a$ and $b$ whenever $a$ is the closest neighbor to $a$ or vice versa, or if the two of them have the same closest neighbor. The initial clusters are then the connected components of this graph. In a second step, some of the clusters get merged in a similar fashion as in hierarchical clustering algorithms until only a previously estimated number of clusters is left. The persistence-based clustering algorithm from from the Topology ToolKit, on the other hand, is similar to AuToMATo insofar as that it produces a persistence diagram of an estimated density function, and then determines how many points in that diagram correspond to “significant" topological features, thus automating the selection of numbers of clusters in ToMATo. Unlike AuToMATo, however, that algorithm does so by making the assumption that the lifetimes of non-significant features roughly follow some exponential, and then declares a point of the estimated persistence diagram to be a feature if it deviates from the estimated exponential curve.
One way of evaluating the clustering performance of an algorithm is to compare the clusterings found by the algorithm to a ground truth clustering. There are many metrics that may be used to that end: for example, set-matching-based, information theoretic, and pair-counting-based metrics [RF16, VEB10]. Set-matching-based metrics rely on finding matches between clusters in the two clusterings. A notable representative of this is the classification error rate, which is often employed in supervised learning. Information-theoretic-based metrics are built upon fundamental concepts from information theory and include the mutual information (MI) score [VEB10], the normalized mutual information score (NMI) etc. Pair-counting-based measures are built upon counting pairs of items on which two clusterings agree or do not agree. Examples of measures that fall into this category are the Rand index [Ran71] and the Fowlkes-Mallows score [FM83], which we use in this paper.
The Fowlkes-Mallows score was originally defined for hierarchical clusterings only, however, it may be defined for general clusterings as follows. Given a clustering $C$ found by an algorithm and a ground truth clustering $G$, one defines the Fowlkes-Mallows score as
$\mathrm{FMS}:=\sqrt{\frac{\mathrm{TP}}{\mathrm{TP}+\mathrm{FP}}}\cdot\sqrt{\frac{\mathrm{TP}}{\mathrm{TP}+\mathrm{FN}}},$
where
In other words, the Fowlkes-Mallows score is defined as the geometric mean of precision and recall of a classifier whose relevant elements are pairs of points that belong to the same cluster in both $C$ and $G$. As such, this score may attain any value between 0 and 1, and these extremal values correspond to the worst and best possible clustering, respectively. We chose to use the Fowlkes-Mallows score as opposed to e.g. mutual information or any of the Rand indices, because the latter have been shown to exhibit biased behaviour depending on whether the clusters in the ground truth are mostly of similar sizes or not, see e.g. [RNBV16]; to the best of our knowledge, the Fowlkes-Mallows score does not suffer from such drawbacks. Moreover, we chose not to use any intrinsic measures of clustering performance since any such measure implicitly defines a further clustering algorithm to compare AuToMATo against, whereas we are interested in comparing AuToMATo against a predefined ground truth clustering.
@@ 3.1 Methodology of AuToMATo
AuToMATo builds upon the ToMATo clustering scheme introduced in [CGOS13] and implemented in [Gli23]. AuToMATo automates the step of visual inspection of the persistence diagram by means of the bottleneck bootstrap, thus promoting ToMATo to a clustering scheme that does not rely on human input.
@@
We point out that our implementation of AuToMATo comes with default values for each of the objects. Each of these values can, of course, be adjusted by the user. For details on these default values, see Section 3.2.
In the present context the bottleneck bootstrap proceeds as follows. AuToMATo generates $B$ bootstrap subsamples $X_{1}^{*},\dots,X_{B}^{*}$ of $X$, each of the same cardinality as $X$. Then, the underlying ToMATo instance with $\tau=+\infty$ and its neighborhood graph and density function estimators is used to compute the persistence diagram for $X$ and each of $X_{1}^{*},\dots,X_{B}^{*}$, yielding persistence diagrams $\widehat{\mathcal{D}}$ and $\widehat{\mathcal{D}}_{1}^{*},\dots,\widehat{\mathcal{D}}_{B}^{*}$, respectively. Using the bootstrapped diagrams $\widehat{\mathcal{D}}_{1}^{*},\dots,\widehat{\mathcal{D}}_{B}^{*}$, a bottleneck bootstrap is performed on $\widehat{\mathcal{D}}$. This yields a value $\widehat{q}_{\alpha}$ that (asymptotically as $n\to\infty$) satisfies
@@
where $\mathcal{D}$ denotes the persistence diagram of the true, unknown density function from which $X$ was sampled. Thus, points of $\widehat{\mathcal{D}}$ of prominence at least $2\cdot\widehat{q}_{\alpha}/\sqrt{n}$ are declared to be significant features of $\widehat{\mathcal{D}}$, and AuToMATo outputs its underlying ToMATo instance with prominence threshold set to $\tau=2\cdot\widehat{q}_{\alpha}/\sqrt{n}$.
When computing the values $\sqrt{n}W_{\infty}(\widehat{\mathcal{D}}_{i}^{*},\widehat{\mathcal{D}})$, $i=1,\dots,B$, in the bottleneck bootstrap, we only consider points in $\widehat{\mathcal{D}}_{i}^{*}$ and $\widehat{\mathcal{D}}$ with finite lifetimes. The reason for this choice is that we consider peaks with infinite lifetime to be significant a priori. Moreover, some of the bootstrapped diagrams among the $\widehat{\mathcal{D}}_{1}^{*},\dots,\widehat{\mathcal{D}}_{B}^{*}$ have a different number of points with infinite lifetime than the reference diagram $\widehat{\mathcal{D}}$. In these cases, the bottleneck distance of the bootstrapped diagram to the reference diagram is infinite, which heavily distorts the distribution $\widetilde{F}(z)$. This choice is justified by experiments.
@@ 3.2 Implementation of AuToMATo
We implemented AuToMATo in Python, and all of the code with documentation is available on GitHub.333github.com/m-a-huber/AuToMATo For a description of AuToMATo in pseudocode, see Algorithm 1. The algorithm has a worst-case complexity of $O(B(nd+n\log(n)+N^{1.5}\log N))$, where $d$ is the dimensionality of the data and $N$ is the maximal number of off-diagonal points across all relevant persistence diagram (which is generally much smaller than $n$); see Appendix A.1 for details. Note that the factor of $B$ can be significantly decreased through parallelization.
@@
Our implementation of AuToMATo is such that the user can directly pass parameters to the underlying ToMATo instance. If no such arguments are provided AuToMATo uses the default choices for those parameters, as determined by the implementation of ToMATo given in [Gli23]. In particular, AuToMATo uses the $k$-nearest neighbor graph and the (logarithm of the) distance-to-measure density estimators by default, each with $k=10$. Of course, the persistence diagrams produced by ToMATo, and hence the output of AuToMATo, depend on this choice. This can lead to suboptimal clustering performance of AuToMATo; see Section 6.
@@
By default, AuToMATo performs the bootstrap on $B=1000$ subsamples of the input point cloud, and sets the confidence level to $\alpha=0.35$. The choice of this latter parameter means that AuToMATo determines merely a 65% confidence region for the persistence diagram produced by the underlying ToMATo instance. While in bootstrapping the confidence level is often set to e.g. $\alpha=0.05$, the seemingly strange choice of $\alpha=0.35$ in the setting of AuToMATo is justified by experiments. The value of 65% seems to be low enough to offset some of the negative influence of using possibly non-optimized neighborhood graph and density estimators discussed in Section 6, while at the same time being high enough to yield good results when these estimators are chosen suitably. We point out that the value $\alpha=0.35$ (as well as the value $B=1000$) was decided on after running an early implementation of AuToMATo on just a few synthetic data sets. In particular, the choice was made before conducting the experiments in Section 4. AuToMATo is implemented in such a way that the parameter $\alpha$ can be adjusted after fitting and the clustering is automatically updated.
@@ 4.1 Choice of clustering algorithms for comparison
  • the FINCH clustering algorithm [SSS19]; and
  • a clustering algorithm building on ToMATo stemming from the Topology ToolKit (TTK) suite [TFL+18]; in the following, we will refer to this as the TTK-algorithm. [^4] For the Topology ToolKit, see topology-tool-kit.github.io/ (BSD license).
For DBSCAN, HDBSCAN and the hierarchical clustering algorithms mentioned above, we worked with their implementations in scikit-learn.555scikit-learn.org/stable/modules/clustering.html For the FINCH clustering algorithm, we worked with the version available on GitHub.666github.com/ssarfraz/FINCH-Clustering (CC BY-NC-SA 4.0 license) Indeed, we subclassed that version in order to make it compatible with the scikit-learn API. Similarly, we created a scikit-learn compatible version of the TTK-algorithm by combining code from TTK with the description of the algorithm given in [CSTL21, Section 5.2]. While we included DBSCAN and HDBSCAN among the clustering algorithms to compare AuToMATo against because they are standard choices, we chose to include the hierarchical clustering algorithms because they are readily available through scikit-learn. Finally, we chose to include FINCH and the TTK-algorithm because, like AuToMATo, they are out-of-the-box (indeed, parameter-free) methods and are thus especially interesting to compare AuToMATo against.
@@ 4.2 Choice of data sets
The data sets on which we ran AuToMATo and the above clustering algorithms stem from the Clustering Benchmarks suite.777clustering-benchmarks.gagolewski.com/ (CC BY-NC-ND 4.0 license) We chose this collection as it comes with a large variety of different data sets, all of which are labeled by one or more ground truths, allowing for a fair and extensive comparison. The collection contains five recommended batteries of data sets from which we selected those (data set, ground truth)-pairs that we deemed reasonable for a general purpose out-of-the-box clustering algorithm. For instance, we chose not to include the data set named olympic that is part of the wut-battery, but we did include the data set named windows from the same battery. Both of these are illustrated in Figure 3.
We made this choice because AuToMATo is a clustering algorithm that determines clusters depending on connectivity, and topologically speaking, there is only one connected component in the olympic data set.
Finally, we excluded all instances where the ground truth contains data points that are labeled as outliers, as outliers creation is currently an experimental feature in AuToMATo.
@@ 4.3 Methodology of the experiments
We min-max scaled each data set, fitted the clustering algorithms to them, and recorded the clustering performance of each result by computing the Fowlkes-Mallows score of the clustering obtained and the respective ground truth (see Section 2.4 for details on this metric and on our choice to work with it).
Since AuToMATo contains a randomized component (namely, the creation of bootstrap samples), we chose to run it ten times on each data set and report its average performance. Moreover, we set the hyperparameters of the HDBSCAN, FINCH and the TTK-algorithm to their default values (as per their respective implementations). In contrast to this, we let the distance threshold parameter for the DBSCAN and the hierarchical clustering algorithms vary from 0.05 to 1.00 in increments of 0.05, with the goal of comparing AuToMATo against the best and worst performing version of these clustering algorithms.
While we restricted ourselves to instances where the ground truth does not contain any points labeled as outliers, some of the clustering algorithms in our list (DBSCAN and HDBSCAN) label some data points as outliers. In order to prevent these algorithms from getting systematically low Fowlkes-Mallows scores because of these outliers, we removed all the points labeled as outliers by these algorithms, and only computed the Fowlkes-Mallows score on the remaining points, both for these clustering algorithms and for AuToMATo. This of course gives an advantage to DBSCAN and HDBSCAN over AuToMATo. Finally, since AuToMATo is not entirely deterministic, we ran AuToMATo on each data set ten times, and recorded the mean of the Fowlkes-Mallows score.
In order to allow reproducibility, we chose a fixed seed for all our experiments, which can be found in our code on GitHub. We ran our experiments on a laptop with a 12th Gen Intel Core i7-1260P processor running at 2.10GHz.

### v3

@@ 2.2 The bottleneck bootstrap
$\widehat{q}_{\alpha}\coloneqq\inf\{z\mid\widetilde{F}(z)\geq 1-\alpha\}$
to be our estimate of $q_{\alpha}$. This estimate is asymptotically consistent by Theorem 2.4, that is, $\widehat{q}_{\alpha}\xrightarrow{n\to\infty}q_{\alpha}$.
In conclusion, the true, unknown persistence diagram $\mathcal{D}$ is at bottleneck distance of at most $\widehat{q}_{\alpha}/\sqrt{n}$ from $\widehat{\mathcal{D}}$ with probability at least $1-\alpha$. Hence, points of $\widehat{\mathcal{D}}$ that are at $L^{\infty}$-distance at most $\widehat{q}_{\alpha}/\sqrt{n}$ from the diagonal could be matched to the diagonal under the bottleneck distance, and thus a point of $\widehat{\mathcal{D}}$ is declared to be a significant feature iff it is at $L^{\infty}$-distance of at least $\widehat{q}_{\alpha}/\sqrt{n}$ to the diagonal, that is, iff its prominence is at least $2\cdot\widehat{q}_{\alpha}/\sqrt{n}$.
@@ 3.1 Methodology of AuToMATo
AuToMATo builds upon the ToMATo clustering scheme introduced in Chazal et al. (2013) and implemented in Glisse (2025). AuToMATo automates the step of visual inspection of the persistence diagram by means of the bottleneck bootstrap, thus promoting ToMATo to a clustering scheme that does not rely on human input.
@@
We point out that our implementation of AuToMATo comes with default values for each of the objects. Each of these values can, of course, be adjusted by the user. For details on these default values, see Subsection 3.2.
To apply the bottleneck bootstrap as described in Section 2.2, AuToMATo generates $B$ bootstrap subsamples $X_{1}^{*},\dots,X_{B}^{*}$ of $X$, each of the same cardinality as $X$, where $X$ is the data set whose points are to be clustered. Then the underlying ToMATo instance with $\tau=+\infty$ and its neighborhood graph and density function estimators is used to compute the persistence diagram for $X$ and each of $X_{1}^{*},\dots,X_{B}^{*}$, yielding persistence diagrams $\widehat{\mathcal{D}}$ and $\widehat{\mathcal{D}}_{1}^{*},\dots,\widehat{\mathcal{D}}_{B}^{*}$, respectively. Using the bootstrapped diagrams $\widehat{\mathcal{D}}_{1}^{*},\dots,\widehat{\mathcal{D}}_{B}^{*}$, a bottleneck bootstrap is performed on $\widehat{\mathcal{D}}$. This yields a value $\widehat{q}_{\alpha}$ that (asymptotically as $n\to\infty$) satisfies
@@
where $\mathcal{D}$ denotes the persistence diagram of the true, unknown density function from which $X$ was sampled. Thus, points of $\widehat{\mathcal{D}}$ of prominence at least $2\cdot\widehat{q}_{\alpha}/\sqrt{n}$ are declared to be significant features of $\widehat{\mathcal{D}}$, and AuToMATo outputs its underlying ToMATo instance with prominence threshold set to $\tau=2\cdot\widehat{q}_{\alpha}/\sqrt{n}$. This procedure is schematically depicted in Figure 2.
Data set $X$
Bootstrap samples
$X_{1}^{*},\dots,X_{B}^{*}$
ToMATo PD
ToMATo PDs
ToMATo PD with threshold
Figure 2: Schematic of the methodology of AuToMATo: from a data set $X$, the usual ToMATo persistence diagram (with $\tau=+\infty$) is computed. Additionally, the analogous persistence diagrams are computed for the bootstrap samples $X_{1}^{*},\dots,X_{B}^{*}$, which are created from $X$ by drawing with replacement. Finally, the bootstrap procedure (indicated by $\otimes$) is used to compute a prominence threshold for the original persistence diagram.
[…]
@@ 3.2 Implementation of AuToMATo
We implemented AuToMATo in Python, and all code with documentation is publicly available. [^3] The code is archived on Zenodo (doi.org/10.5281/zenodo.17279741) and developed openly on GitHub (github.com/m-a-huber/automato_paper). For a description of AuToMATo in pseudocode, see Algorithm 1. The algorithm has a worst-case complexity of $O(B(nd+n\log(n)+N^{1.5}\log N))$, where $d$ is the dimensionality of the data and $N$ is the maximal number of off-diagonal points across all relevant persistence diagrams (which is generally much smaller than $n$); see below for details. Note that the factor of $B$ can be significantly decreased through parallelization.
@@
Our implementation of AuToMATo is such that the user can directly pass parameters to the underlying ToMATo instance. If no such arguments are provided AuToMATo uses the default choices for those parameters, as determined by the implementation of ToMATo given in Glisse (2025). In particular, AuToMATo uses the $k$-nearest neighbor graph and the (logarithm of the) distance-to-measure density estimators by default, each with $k=10$. Of course, the persistence diagrams produced by ToMATo, and hence the output of AuToMATo, depend on this choice. This can lead to suboptimal clustering performance of AuToMATo; see Section 6.
@@
[…]
Complexity analysis of Algorithm 1:
Recall from Chazal et al. (2013, Section 2) that, if an estimated density and a neighborhood graph are provided, ToMATo has a worst-case time complexity in $O(n\log(n)+m\alpha(n))$, where $n$ and $m$ are the number of vertices and edges of the neighborhood graph, respectively, and $\alpha$ denotes the inverse Ackermann function (note that $n$ equals the number of data points). By default, ToMATo (and hence AuToMATo) works with the $k$-nearest neighbor graph and distance-to-measure density estimators, where the latter itself relies on the $k$-nearest neighbor graph (each with $k=10$). Taking into account the known complexity bound $O(nd)$ for the creation of the $k$-nearest neighbor graph (where $d$ is the dimensionality of the data), and using the fact that $m\in O(n)$ for this graph, this leads to a worst-case time complexity in $O(nd+n\log(n))$ for a single run of ToMATo. Creating the bootstrap samples $X_{i}^{*}$, $i=1,\dots,B$, has complexity in $O(Bn)$; computing the values $\sqrt{n}W_{\infty}^{\mathrm{fin}}(\mathcal{D}_{i}^{*},\mathcal{D})$, $i=1,\dots,B$, has worst-case complexity $O(BN^{1.5}\log(N))$ (where $N$ denotes the maximal number of off-diagonal points across all relevant persistence diagrams), and sorting them has worst-case complexity in $O(B\log(B))$. Combined, this leads to a worst-case complexity for AuToMATo in $O(B(nd+n\log(n)+N^{1.5}\log(N))+B\log(B))$. Using that $B$ is a constant, we obtain the runtime of $O(B(nd+n\log(n)+N^{1.5}\log N))$ claimed above.
We point out that the complexity of $O(N^{1.5}\log(N))$ for the computation of the bottleneck distance between a pair of persistence diagrams is a theoretical worst-case scenario, whose validity was established in Efrat et al. (2001). In practice, one typically allows for a small error in the computation of bottleneck distances which decreases the effective complexity (Kerber et al., 2017). Our implementation of AuToMATo makes use of this by allowing for the smallest possible error as determined by the smallest positive double (see the implementation in  Godi (2025) for details). Moreover, note that the factor $B$ appearing in the complexity of AuToMATo can be drastically decreased in practice through parallelization.
The Python package:
@@ 4.1 Choice of clustering algorithms for comparison
  • the FINCH clustering algorithm (Sarfraz et al., 2019); and
  • a clustering algorithm building on ToMATo stemming from the Topology ToolKit (TTK) suite (Tierny et al., 2018); in the following, we will refer to this as the TTK-algorithm. [^4] For the Topology ToolKit, see topology-tool-kit.github.io/ (BSD license).
For DBSCAN, HDBSCAN and the hierarchical clustering algorithms mentioned above, we worked with their implementations in scikit-learn.555scikit-learn.org/stable/modules/clustering.html For the FINCH clustering algorithm, we worked with the version available on GitHub.666github.com/ssarfraz/FINCH-Clustering (CC BY-NC-SA 4.0 license) Indeed, we subclassed that version in order to make it compatible with the scikit-learn API. Similarly, we created a scikit-learn compatible version of the TTK-algorithm by combining code from TTK with the description of the algorithm given in Cotsakis et al. (2021, Section 5.2). While we included DBSCAN and HDBSCAN among the clustering algorithms to compare AuToMATo against because they are standard choices, we chose to include the hierarchical clustering algorithms because they are readily available through scikit-learn. Finally, we chose to include FINCH and the TTK-algorithm because, like AuToMATo, they are out-of-the-box (indeed, parameter-free) methods and are thus especially interesting to compare AuToMATo against.
@@ 4.2 Choice of data sets
The data sets on which we ran AuToMATo and the above clustering algorithms stem from the Clustering Benchmarks suite (Gagolewski, 2022). [^7] Specifically, we worked with version 1.1.0 of the benchmarking suite (Gagolewski et al., 2022) We chose this collection as it comes with a large variety of different data sets, all of which are labeled by one or more ground truths, allowing for a fair and extensive comparison. The collection contains five recommended batteries of data sets from which we selected those (data set, ground truth)-pairs that we deemed reasonable for a general purpose parameter-free clustering algorithm. For instance, we chose to include the data set named windows that is part of the wut-battery, but not the data set named windows from the same battery (see Figure 6 in the appendix for an illustration). We chose to include the windows data set because AuToMATo determines clusters depending on connectivity, and topologically speaking, there is only one connected component in the olympic data set. Finally, we excluded all instances where the ground truth contains data points that are labeled as outliers, as outliers creation is currently an experimental feature in AuToMATo.
@@ 4.3 Methodology of the experiments
We min-max scaled each data set, fitted the clustering algorithms to them, and recorded the clustering performance of each result by computing the Fowlkes-Mallows score (Fowlkes & Mallows, 1983) of the clustering obtained and the respective ground truth. While the Fowlkes-Mallows score was originally defined for hierarchical clusterings only, it may be defined for general clusterings as follows. Given a clustering $C$ found by an algorithm and a ground truth clustering $G$, one defines the Fowlkes-Mallows score as
$\mathrm{FMS}\coloneqq\sqrt{\frac{\mathrm{TP}}{\mathrm{TP}+\mathrm{FP}}}\cdot\sqrt{\frac{\mathrm{TP}}{\mathrm{TP}+\mathrm{FN}}},$
where
In other words, the Fowlkes-Mallows score is defined as the geometric mean of precision and recall of a classifier whose relevant elements are pairs of points that belong to the same cluster in both $C$ and $G$. It may attain any value between 0 and 1, and these extremal values correspond to the worst and best possible clustering, respectively. We chose to use the Fowlkes-Mallows score as opposed to, for instance, mutual information or any of the Rand indices, because the latter have been shown to exhibit biased behavior depending on whether the clusters in the ground truth are mostly of similar sizes or not, see, for instance, Romano et al. (2016); to the best of our knowledge, the Fowlkes-Mallows score does not suffer from such drawbacks. Moreover, we chose not to use any intrinsic measures of clustering performance since any such measure implicitly defines a further clustering algorithm to compare AuToMATo against, whereas we are interested in comparing AuToMATo against a predefined ground truth clustering.
We set the hyperparameters of the HDBSCAN, FINCH and the TTK-algorithm to their default values (as per their respective implementations). In contrast to this, we let the distance threshold parameter for the DBSCAN and the hierarchical clustering algorithms vary from 0.05 to 1.00 in increments of 0.05, with the goal of comparing AuToMATo against the best and worst performing version of these clustering algorithms. To account for the randomized component of AuToMATo, we ran it ten times, each time with a different seed.
While we restricted ourselves to instances where the ground truth does not contain any points labeled as outliers, some of the clustering algorithms in our list (DBSCAN and HDBSCAN) label some data points as outliers. In order to prevent these algorithms from getting systematically low Fowlkes-Mallows scores because of these outliers, we removed all the points labeled as outliers by these algorithms, and only computed the Fowlkes-Mallows score on the remaining points, both for these clustering algorithms and for AuToMATo. This of course gives an advantage to DBSCAN and HDBSCAN over AuToMATo.
In order to allow reproducibility, we chose a fixed seed for all our experiments, which can be found in our code. [^8] The code is archived on Zenodo (doi.org/10.5281/zenodo.17279741) and developed openly on GitHub (github.com/m-a-huber/automato_paper). We ran our experiments on a laptop with a 12th Gen Intel Core i7-1260P processor running at 2.10GHz.